# MoESD: Unveil Speculative Decoding’s Potential for Accelerating Sparse MoE

### v2

@@ 1 Introduction
Speculative decoding (SD) is a lossless technique to accelerate LLM inference, but conventional wisdom suggests that its efficacy diminishes when applied to MoEs. In SD, a smaller draft model is introduced to rapidly generate multiple candidate tokens, while the larger target model verifies these predictions in parallel, preserving only correctly speculated tokens. For dense models’ inference, the time taken to generate a single token and verify multiple ones is roughly the same, as both tasks require the full set of parameters to be loaded once. Therefore, SD gains acceleration through fewer forward rounds of the target model and shorter decoding time of the draft model. However, this acceleration is considered to diminish in MoEs li2024eagle , as the multiple draft tokens in verification activate more experts than a single token, leading to larger memory access and significantly longer verification time compared to a standard decoding step.
@@
The insight above is supported by comprehensive theoretical analyses, through which we identify a new metric target efficiency to quantify how systemic factors (such as workload and target model architecture) affect the speedup by SD. In contrast to the acceptance rate, which is an algorithmic metric to evaluate SD in existing works chen2023accelerating ; leviathan2023fast ; li2024eagle ; cai2401medusa , the proposed target efficiency isolates the extrinsic factors of speculating algorithm selection, enabling researchers to focus on the intrinsic system bottlenecks caused by the target model’s computational and memory access requirements. As a further step, we build a quantitative model of SD speedup for MoE based on these analyses. Our modeling is validated to be reliable through comparison with experimental results. Besides, it provides an approach for analyzing the execution time of different components, making the end-to-end SD acceleration results more transparent and explainable.
Our work offers a new perspective for lossless MoE acceleration, particularly well-suited for private serving scenarios privatellm1 ; privatellm2 ; privatellm3 . Private serving has gained popularity among companies seeking to safeguard their data and model security, as it restricts access to organizational members only. These environments typically process moderate batches containing tens of requests. SD can also effectively accelerate memory-restricted scenarios where MoEs exceed GPU capacity.
@@
  • Based on theoretical analysis, we developed a reliable modeling for SD speedup, thus making the acceleration process transparent and explainable. Existing metrics only assess algorithmic optimization efficiency and cannot fully explain SD speedup, so we introduce a new systematic metric target efficiency that reveals speedup opportunities inherent in the target model.
@@ 2 Related Work
Speculative decoding (SD), initially proposed by leviathan2023fast and chen2023accelerating , has emerged as a widely adopted technique for accelerating LLM inference without sacrificing generation quality. In its most fundamental form, SD employs a smaller model to rapidly generate draft tokens, which are then verified in parallel by the target model that needs to be accelerated. Subsequently, more algorithms are developed to lift the acceptance rate of draft tokens. miao2024specinfer ; cai2401medusa ; he2023rest ; svirschevski2024specexec ; li2024eagle ; li2024eagle2 ; li2025eagle3 adopt tree-structured generation patterns rather than chains to explore a broader range of potential completions. liu2024deepseek ; cai2401medusa ; li2024eagle ; li2024eagle2 ; li2025eagle3 propose to replace draft models with specifically trained speculative heads integrated in the target model.
Despite advances in SD algorithms, it has long been considered ineffective for large batches or MoE li2024eagle , since the verification time in these cases significantly increases. Until recently, MagicDec sadhukhan2024magicdec first challenged that in long-sequence regimes, SD can effectively accelerate large batches, primarily due to the significantly increased KV cache altering the computation-to-memory access ratio of the model. However, SD research for MoE remains unexplored. In response, our work fills this gap, unveiling that under certain conditions, SD can effectively accelerate MoE models. Since our focus is not on algorithmic optimization, this paper employs the most fundamental chain-like speculation pattern.
@@ 3 Theoretical Analysis
In this section, we present the theoretical analyses supporting our conclusion that SD can be more effective for MoE than dense models at moderate batch sizes. We begin by formalizing general SD speedup and introducing our new metric target efficiency (Sec. 3.1). Then, we focus on MoEs, analyzing how workload and MoE sparsity collectively affect the number of activated experts and SD speedup (Sec. 3.2). Based on these analyses, we develop a performance model that aligns with GPU results (Sec. 3.3). We further discuss the application scenarios of our theoretical findings (Sec. 3.4).
@@
RP=peak computation power (unit: Flops)peak memory bandwidth (unit: bytes/second)AI=computation operation (unit: times)memory access volume (unit: $\small\textbf{RP}=\frac{\text{peak computation power {(unit: Flops)}}}{\text{peak memory bandwidth {(unit: bytes/second)}}}\quad\textbf{AI}=\frac{\text{computation operation {(unit: times)}}}{\text{memory access volume {(unit: bytes)}}}$ (1)
@@ 3.1 Formulation of Speculative Decoding Speedup and Target Efficiency
We first formalize the processing time of speculative decoding, denoted as $T_{SD}$. For generating a sequence of length $S$, SD goes through $R$ rounds, each containing three stages: ① the draft model proposes $\gamma$ tokens as specified by the speculation strategy; ② the target model verifies these tokens; ③ rejection sampling chen2023accelerating discards incorrectly predicted tokens based on logits from target and draft models. We use $T_{T}(b,s)$ and $T_{D}(b,s)$ to represent the time for once forwarding of the target and draft model, respectively, where $b$ and $s$ are the formal arguments for batch size and the number of tokens to process. [^1] Since we work with typical sequence lengths and moderate batch sizes, the impact of KV-cache on performance is limited, allowing us to omit the already generated sequence length from our analysis. For cases where KV-cache becomes the dominant factor, see sadhukhan2024magicdec . Therefore, the time for processing a batch containing $B$ requests is given by:
$T_{SD}=R\times(T_{propose}+T_{verify}+T_{reject})=R\times\Big{(}\gamma\cdot T_{D}(B,1)+T_{T}(B,\gamma)+T_{reject}\Big{)}\$ (2)
@@
The denominator of Eq. 4 consists of three terms. $\frac{T_{D}(B,1)}{T_{T}(B,1)}$ is the ratio of draft-model forward time over target-model forward time, reflecting the relative volume of draft and target models. It is also kept small (usually less than 1/10) to ensure the speculation is efficient. $\frac{T_{reject}}{T_{T}(B,1)}$ is even smaller, since $T_{reject}$ only involves sampling rather than model inference. $\frac{T_{T}(B,\gamma)}{T_{T}(B,1)}$, which is the ratio of multi-token forward time over single-token forward time, has the biggest value among these three items and significantly affects the final speedup. As indicated by Eq. 4, its increase causes speedup reduction. Two different factors drive its increase, explaining SD’s ineffectiveness under large batches with both dense models and MoE, and under small batches with MoE, respectively:
  1. 1.
    The compute-boundness. The model’s $\frac{T_{T}(B,\gamma)}{T_{T}(B,1)}$ approaches 1 when more memory-bound (smaller $B$s) but increases to $\gamma$ when more compute-bound (larger $B$s).
  2. 2.
@@
Therefore, we define target efficiency as $\frac{T_{T}(B,1)}{T_{T}(B,\gamma)}$, which helps understand the underlying causes of SD acceleration variation across different scenarios. Our experiments shown in Fig. 2 demonstrate that target efficiency consistently reflects the trend of SD speedup variations. Despite the importance of this value, previous works have rarely noticed it, primarily due to differences in research focus. Previous SD research mainly addresses the question by lifting the acceptance rate:
@@ 3.2 Moderate Batch Size Enables Speculative Decoding Speedup for MoE
We use the Bernoulli random variable $X$ to indicate the activation for experts: $X_{i}=1$ for expert $i$ being activated, 00 otherwise. For simplicity, we assume $X$s are i.i.d. Then, the expected number of activated experts $N$ can be expressed as Eq. 6, where $E$ denotes the total expert count and $Pr(X_{i})$ represents the probability that the $i^{th}$ expert is activated.
@@
$Pr(X)=1-Pr(\text{ None of the $t$ tokens activates the expert })=1-(\frac{E-K}{E})^{t}$ (7)
@@
Therefore, when $B$ exceeds $T_{thres}$, the number of activated experts saturates, causing the $B\gamma$ tokens in verification to incur only marginally larger memory access. Having addressed the second factor for $\frac{T_{T}(B,\gamma)}{T_{T}(B,1)}$’s increase analyzed in Sec. 3.1, we now turn to the potential compute limitations caused by the first factor. If such $B$s make the system compute-bound, SD would also fail to accelerate MoE effectively. Our answer to this concern is: Sparser MoEs delay the transition from memory-bound to compute-bound when input tokens count increases.
@@
As proven in Appendix A and shown in Figure 1(c), given $t=T>1$, $\overline{T_{exp}}(T;\rho)$ decreases with $\rho$, indicating that as MoE becomes sparser, each expert processes fewer tokens per parameter loading. Consequently, the system running sparser MoEs is more memory-bound, leading to lower utilization of arithmetic units. The verification stage can therefore leverage these spare resources without notably increasing processing time. In contrast, dense models are extreme cases with $\rho=1$, where the FFN consistently approaches the maximal possible arithmetic intensity of $T$, and the system transitions rapidly to the compute-bound regime as $T$ increases.
@@ 3.3 A Modeling Method for Speculative Decoding Speedup
4:      bias, $k_{1}$, $k_{2}$, $k_{3}$, draft_bias, draft_k, reject_bias, reject_k, $\lambda$, s $=$ params $\triangleright$ Unpack parameters
5:      $N_{ar}=E\cdot(1-((E-K)/E)^{B}),~{}~{}T_{ar}=B\cdot K/N_{ar}$triangleright Compute AR forward time
6:      ar_time $=$ bias $+k_{1}\cdot G(B;\lambda RP,s)+k_{2}\cdot N_{ar}+k_{3}\cdot G(T_{ar};\lambda RP,s)$
7:      $N_{sd}=E\cdot(1-((E-K)/E)^{B\gamma}),~{}~{}T_{sd}=B\cdot\gamma\cdot K/N_{sd}$triangleright Compute SD forward time
8:      verify_time $=$ bias $+k_{1}\cdot G(B\gamma;\lambda RP,s)+k_{2}\cdot N_{sd}+k_{3}\cdot G(T_{sd};\lambda RP,s)$
9:      draft_time $=$ draft_bias + draft_k $\cdot~{}G(B;\lambda RP,s)$triangleright Compute draft model forward time
10:      reject_time $=$ reject_bias + reject_k $\cdot B$triangleright Compute rejection sampling time
11:      Speedup $=\sigma\cdot(\gamma+1)\cdot\frac{\textit{\text{ar\_time}}}{\textit{\text{draft\_time}}+\textit{\text{ar\_time}}+\textit{\text{verify\_time}}+\text{reject\_time}}$triangleright Compute the speedup as formalized in Eq. 4
@@
For the MoE target model, factors (1), (2), and (3) are all involved. We combine them in a first-order style and introduce parameters bias, $k_{1}$, $k_{2}$, and $k_{3}$ to adjust for non-ideal factors in actual GPU execution, with the full expression shown in lines 6 and 8 of Alg. 1. These parameters have clear practical meanings: bias represents the time required to load fixed parameters; $k_{2}\cdot N$ represents the time needed to load $N$ activated experts; $k_{1}\cdot G(t)$ and $k_{3}\cdot G(\overline{T_{exp}})$ describe the incremental trend in execution time as the number of tokens increases. For the draft model, only factor (1) is involved since it is usually dense, with the modeling form shown in line 9 of Alg. 1. Afterwards, we leverage the measurement inputs to automatically determine the optimal parameter values, with the optimization criterion being the minimization of Mean Squared Error (MSE) between the model outputs and the ground truth, as shown in line 13 of Alg. 1. Applying these optimized parameters in our model (i.e., the ComputeSpeedup function in line 3), we can then obtain the SD speedup for any given workload.
@@
3.4 From Theory to Practice: Application Scenarios
While the previous section focused on theoretical analysis, here we demonstrate how these findings translate to practical speedups. SD provides the greatest benefits for MoE models with moderate batch sizes, making it particularly suitable for private serving scenarios that fall between edge and cloud deployments. These scenarios actually face an ’efficiency gap’ in MoE: they are forced to load all parameters (unlike the edge that selectively activate experts), yet don’t have sufficient computation to fully utilize GPU FLOPs (unlike compute-bound public cloud senarios). In turn, our finding offers a novel solution to address this efficiency challenge without compromising model quality.
When MoE models are too large to fit entirely in GPU memory and offloading becomes necessary, SD can achieve acceleration across a wider range of batch sizes. FFNs typically comprise over 90% of total MoE parameters, so usually experts are offloaded to CPU memory while other parameters are kept on GPU, as implemented in k-Transformers ktransformers . The parameter loading bandwidth then degrades from GPU memory bandwidth to the much lower PCIe bandwidth, thus making the system more memory-bound and leaving more computational resources underutilized. Consequently, hardware can process additional tokens without significantly increasing processing time, creating favorable conditions for SD. In contrast, under moderate batch sizes, existing optimizations like expert prefetching and caching lose efficiency since nearly all experts are activated.
Most analyses in this paper do not specifically consider expert parallelism (EP), but they are compatible with EP. While EP distributes experts across multiple GPUs, it affects neither the number of activated experts $N(t)$ nor the average number of tokens per expert $\overline{T_{exp}}(t;\rho)$ that we previously derived. Given that current state-of-the-art MoE models are equipped with a large number of experts, each EP device typically holds multiple experts. Consequently, on each device, multi-token verification is expected to activate more experts than the single-token decoding, consistent with our earlier analysis. Since components besides MoE FFN are also parallelized, MoE FFN still constitutes a significant portion of processing time, thus allowing the impact of its memory-boundedness to remain observable in end-to-end performance. Changes happen when EP scales so extensively that verification no longer incurs more expert accesses than decoding on each GPU. For instance, when the number of EP devices equals the expert count, both verification and decoding load only one expert per GPU even at small batch sizes, though verification engages more GPUs for computation of more tokens. Under these conditions, verification becomes only marginally longer than decoding, thereby restoring the efficiency of SD. This improvement is fundamentally enabled by the increased aggregate memory bandwidth provided by the larger number of EP GPUs.

### v3

@@ 1 Introduction
Speculative decoding (SD) is a lossless technique to accelerate LLM inference, but conventional wisdom suggests that its efficacy diminishes when applied to MoEs. In SD, a smaller draft model is introduced to rapidly generate multiple candidate tokens, while the larger target model verifies these predictions in parallel, preserving only correctly speculated tokens. For dense models’ inference, the time taken to generate a single token and verify multiple ones is roughly the same, as both tasks require the full set of parameters to be loaded once. Therefore, SD gains acceleration through fewer forward rounds of the target model and shorter decoding time of the draft model. However, this acceleration has been demonstrated to diminish in MoEs li2024eagle ; moesd-inefficiency , as the multiple draft tokens in verification activate more experts than a single token, leading to larger memory access and significantly longer verification time compared to a standard decoding step.
@@
The insight above is supported by comprehensive theoretical analyses, through which we identify a new metric target efficiency to quantify how systemic factors (such as workload and target model architecture) affect SD speedup. In contrast to existing SD works that use acceptance rate chen2023accelerating ; leviathan2023fast ; li2024eagle ; cai2401medusa , an algorithmic metric to evaluate how accurately the draft model speculates the target model, our proposed target efficiency isolates extrinsic factors like algorithm selection and focuses on intrinsic system bottlenecks caused by the target model’s computational and memory access requirements. As demonstrated in the following sections, even with similar acceptance rates, systemic factors can greatly impact SD effectiveness, making our metric target efficiency necessary for a comprehensive understanding of SD acceleration.
As a further step, we build a quantitative modeling of SD speedup for MoE based on these theoretical analyses. The consistent matching between our modeling and experiment results confirms the reliability of our analyses. Additionally, the modeling itself provides an approach for analyzing the execution time of different components, making the end-to-end SD acceleration results more transparent and explainable.
Our work offers a new perspective for lossless MoE acceleration, particularly well-suited for private serving scenarios privatellm1 ; privatellm2 ; privatellm3 . Private serving has gained popularity among enterprises seeking to safeguard data and model security, with typical applications such as in-house chatbots. These environments typically process moderate batches containing tens of requests. Additionally, our findings can be applied to latency-critical scenarios where large batch sizes are infeasible, or memory-restricted environments where MoEs exceed GPU capacity.
@@
  • Based on theoretical analysis, we developed a reliable modeling for SD speedup, thus making the acceleration process transparent and explainable. Existing metrics only assess algorithmic optimization efficiency and cannot fully explain SD speedup, so we introduce a new systemic metric target efficiency that reveals speedup opportunities inherent in the target model.
@@ 2 Related Work
Speculative decoding (SD), initially proposed by leviathan2023fast and chen2023accelerating , has emerged as a widely adopted technique for accelerating LLM inference without sacrificing generation quality. Basic SD employs a smaller model to rapidly generate draft tokens, which are then verified in parallel by the target model that needs to be accelerated. Afterwards, more algorithms are developed to lift the acceptance rate of draft tokens. miao2024specinfer ; cai2401medusa ; he2023rest ; svirschevski2024specexec ; li2024eagle ; li2024eagle2 ; li2025eagle3 adopt tree-structured generation patterns rather than chains to explore a broader range of potential completions. liu2024deepseek ; cai2401medusa ; li2024eagle ; li2024eagle2 ; li2025eagle3 propose to replace draft models with specifically trained speculative heads integrated in the target model.
Despite advances in SD algorithms, it has long been considered ineffective for large batches batchsd2 ; batchsd3 ; miao2024specinfer or MoE li2024eagle ; moesd-inefficiency , since the verification time in these cases significantly increases. Until recently, MagicDec sadhukhan2024magicdec first challenged that in long-sequence regimes, SD can effectively accelerate large batches, primarily due to the significantly increased KV cache altering the computation-to-memory access ratio of the model. However, SD research for MoE remains unexplored. In response, our work fills this gap, unveiling that under certain conditions, SD can effectively accelerate MoE models.
@@ 3 Theoretical Analysis
In this section, we present the theoretical analyses supporting our conclusion that SD can be more effective for MoE than dense models at moderate batch sizes. We begin by formalizing general SD speedup and introducing our new metric target efficiency (Sec. 3.1). Then, we focus on MoEs, analyzing how workload and MoE sparsity collectively affect the number of activated experts and SD speedup (Sec. 3.2). Based on these analyses, we develop a performance model that aligns with GPU results (Sec. 3.3). We further discuss the practical value of our theoretical findings (Sec. 3.4).
@@
[…]
@@ 3.1 Formulation of Speculative Decoding Speedup and Target Efficiency
We first formalize the processing time of speculative decoding, denoted as $T_{SD}$. To generate a sequence of length $S$, SD goes through $R$ rounds, each containing three stages: ① the draft model proposes $\gamma$ tokens as specified by the speculation strategy; ② the target model verifies these tokens; ③ rejection sampling chen2023accelerating discards incorrectly predicted tokens based on logits from target and draft models. We use $T_{T}(b,s)$ and $T_{D}(b,s)$ to represent the time for once forwarding of the target and draft model, respectively, where $b$ and $s$ are the formal arguments for batch size and the number of tokens to process. [^1] Since we work with typical sequence lengths and moderate batch sizes, the impact of KV-cache on performance is limited, allowing us to omit the already generated sequence length from our analysis. For cases where KV-cache becomes the dominant factor, see sadhukhan2024magicdec . Therefore, the time for processing a batch containing $B$ requests is given by:
$T_{SD}=R\times(T_{propose}+T_{verify}+T_{reject})=R\times\Big(\gamma\cdot T_{D}(B,1)+T_{T}(B,\gamma)+T_{reject}\Big)\\$ (2)
@@
The denominator of Eq. 4 consists of three terms. $\frac{T_{D}(B,1)}{T_{T}(B,1)}$ is the ratio of draft-model forward time over target-model forward time, reflecting the relative volume of draft and target models. It is also kept small (usually less than 1/10 miao2024specinfer ; li2024eagle ; leviathan2023fast ) to ensure the speculation is efficient. $\frac{T_{reject}}{T_{T}(B,1)}$ is even smaller, since $T_{reject}$ only involves sampling rather than model inference. $\frac{T_{T}(B,\gamma)}{T_{T}(B,1)}$, which is the ratio of multi-token forward time over single-token forward time, has the biggest value among these three items and significantly affects the final speedup. As indicated by Eq. 4, its increase causes speedup reduction. Two different factors drive its increase, explaining SD’s ineffectiveness under (1) large batches for both dense models and MoE, and (2) MoE with small batches, respectively:
  1. (1)
    The compute-boundness. The model’s $\frac{T_{T}(B,\gamma)}{T_{T}(B,1)}$ approaches 1 when more memory-bound (smaller batch size $B$) but increases to $\gamma$ when more compute-bound (larger batch size $B$).
  2. (2)
@@
Therefore, we define target efficiency as $\frac{T_{T}(B,1)}{T_{T}(B,\gamma)}$, which helps understand the systemic causes of SD acceleration degradation as listed above. Our experiments shown in Fig. 2 demonstrate that target efficiency consistently reflects the trend of SD speedup variations. Despite the importance of this value, previous works have rarely noticed it, primarily due to differences in research focus. Previous SD research mainly addresses the question by lifting the acceptance rate:
@@ 3.2 Moderate Batch Size Enables Speculative Decoding Speedup for MoE
We use the Bernoulli random variable $X$ to indicate the activation for experts: $X_{i}=1$ for expert $i$ being activated, 0 otherwise. For simplicity, we assume $X$s are i.i.d. Then, the expected number of activated experts $N$ can be expressed as Eq. 6, where $E$ denotes the total expert count and $Pr(X_{i})$ represents the probability that the $i^{th}$ expert is activated.
@@
[…]
@@
Therefore, when $B$ exceeds $T_{thres}$, the number of activated experts saturates, causing the $B\gamma$ tokens in verification to incur only marginally larger memory access. Having addressed the second factor (namely, extra memory loads) for $\frac{T_{T}(B,\gamma)}{T_{T}(B,1)}$’s increase analyzed in Sec. 3.1, we now turn to the potential limitations caused by the first factor of compute-boundness. If such $B$s make the system compute-bound, SD would also fail to accelerate MoE effectively. Our answer to this concern is: Sparser MoEs delay the transition from memory-bound to compute-bound when input tokens count increases.
@@
As proven in Appendix and shown in Figure 1(c), given $t=T>1$, $\overline{T_{exp}}(T;\rho)$ decreases with $\rho$, indicating that as MoE becomes sparser, each expert processes fewer tokens per parameter loading. Consequently, the system running sparser MoEs is more memory-bound, leading to lower utilization of arithmetic units. The verification stage can therefore leverage these spare resources without notably increasing processing time. In contrast, dense models are extreme cases with $\rho=1$, where the FFN consistently approaches the maximal possible arithmetic intensity of $T$, and the system transitions rapidly to the compute-bound regime as $T$ increases.
@@ 3.3 A Modeling Method for Speculative Decoding Speedup
4:   bias, $k_{1}$, $k_{2}$, $k_{3}$, draft_bias, draft_k, reject_bias, reject_k, $\lambda$, s == params $\triangleright$ Unpack parameters
5:   $N_{ar}=E\cdot(1-((E-K)/E)^{B}),~~T_{ar}=B\cdot K/N_{ar}$ $\triangleright$ Compute AR forward time
6:   ar_time == bias $k_{1}\cdot G(B;\lambda RP,s)+k_{2}\cdot N_{ar}+k_{3}\cdot G(T_{ar};\lambda RP,s)$
7:   $N_{sd}=E\cdot(1-((E-K)/E)^{B\gamma}),~~T_{sd}=B\cdot\gamma\cdot K/N_{sd}$ $\triangleright$ Compute SD forward time
8:   verify_time == bias $k_{1}\cdot G(B\gamma;\lambda RP,s)+k_{2}\cdot N_{sd}+k_{3}\cdot G(T_{sd};\lambda RP,s)$
9:   draft_time == draft_bias + draft_k $\cdot~G(B;\lambda RP,s)$ $\triangleright$ Compute draft model forward time
10:   reject_time == reject_bias + reject_k $\cdot B$ $\triangleright$ Compute rejection sampling time
11:   Speedup $\sigma\cdot(\gamma+1)\cdot\frac{\textit{\text{ar\_time}}}{\textit{\text{draft\_time}}+\textit{\text{ar\_time}}+\textit{\text{verify\_time}}+\text{reject\_time}} \triangleright$ Compute the speedup as formalized in Eq. 4
@@
For the MoE target model, factors (1), (2), and (3) are all involved. We combine them in a first-order style and introduce parameters bias, $k_{1}$, $k_{2}$, and $k_{3}$ to adjust for non-ideal factors in actual GPU execution, with the full expression shown in lines 6 and 8 of Alg. 1. These parameters have clear practical meanings: bias represents the time required to load fixed parameters; $k_{2}\cdot N$ represents the time needed to load $N$ activated experts; $k_{1}\cdot G(t)$ and $k_{3}\cdot G(\overline{T_{exp}})$ describe the incremental trend in execution time as the number of tokens increases. For the draft model, only factor (1) is involved since it is usually dense, with the modeling form shown in line 9 of Alg. 1.
With the expression of SD speedup determined, we fit the measurement inputs to automatically determine the relaxation parameter values, with the optimization criterion being the minimization of Mean Squared Error (MSE) between the model outputs and the ground truth, as shown in line 13 of Alg. 1. By applying these optimized parameters in our model (i.e., the ComputeSpeedup function in line 3), we obtain the complete modeling for SD speedup. An illustrative diagram of this process and more fitting details are provided in Appendix C.
@@
3.4 Practical Values of Theoretical Findings
While previous sections focuses on theoretical analysis, this section demonstrates how these findings translate to practical speedups. Our theoretical analysis has already revealed that SD speedup for MoE is most effective at moderate batch sizes, with its trend initially increasing and then decreasing. We discuss their practical values considering both basic deployment and extended configurations.
Basic deployment.
(1) Moderate batch sizes are common in private serving, which are increasingly adopted for data security, with representative applications like enterprise in-house chatbots. (2) When latency requirements are strict, large batch sizes are often not feasible. LLM serving must satisfy multiple service level objectives (SLOs) goodput , including time-to-first-token (TTFT) and time-per-output-token (TPOT). Large batches reduce per-request computational resources, causing latency violations. In such cases, moderate batch sizes are common. (3) Our work actually reveals that SD on MoE relaxes the traditional latency-throughput trade-off. Specifically, MoE models exhibit a regime where SD speedup increases (lower latency) alongside larger batch sizes (higher throughput).
From the model’s perspective, moderate batch sizes represent an "efficiency gap" in MoEs. At this scale, all parameters must be loaded (unlike small batches with selective expert activation), yet GPU FLOPs are not fully utilized (unlike large batches). Our findings provide a novel perspective to address this efficiency challenge without compromising model quality.
Extended configurations.
We consider typical system optimizations on MoE like offloading and expert parallelism (EP). When MoE models exceed GPU memory capacity, FFN parameters are offloaded to CPU memory ktransformers . This degrades parameter loading bandwidth from GPU memory bandwidth to much lower PCIe bandwidth, making the system more memory-bound. Consequently, additional computation does not significantly increase processing time, creating favorable conditions for SD. Notably, existing optimizations like expert prefetching xue2024moe ; zhong2024adapmoe and caching he2024expertflow ; tang2024hobbit lose efficiency under moderate batch sizes since nearly all experts are activated.
Our findings are also compatible with EP. In EP, experts are distributed across multiple GPUs, which affects neither $N(t)$ nor $\overline{T_{exp}}$, making our previous analyses remain valid. Since components besides MoE FFN are also parallelized, MoE FFN continues to constitute a significant portion of processing time, allowing memory-boundness effects to remain observable in end-to-end performance. Notably, under extensive EP configurations, the inefficiency of SD for MoE at a small batch size may vanish, considering the additional memory bandwidth offered by large amounts of EP GPUs.